### v1

# Enhanced many-body quantum scars from the non-Hermitian Fock skin effect

Abstract
In contrast with extended Bloch waves, a single particle can become spatially localized due to the so-called skin effect originating from non-Hermitian pumping. Here we show that in a wide class of kinetically-constrained many-body systems, the skin effect can instead manifest as dynamical amplification within the Fock space, beyond the intuitively expected and previously studied particle localization and clustering. We exemplify this non-Hermitian Fock skin effect in an asymmetric version of the PXP model and show that it gives rise to ergodicity-breaking eigenstates – the non-Hermitian analogs of quantum many-body scars. A distinguishing feature of these non-Hermitian scars is their enhanced robustness against external disorders. We propose an experimental realization of the non-Hermitian scar enhancement in a tilted Bose-Hubbard optical lattice with laser-induced loss. Our results show that the Fock skin effect provides a powerful tool for creating robust non-ergodic states in generic open quantum systems.
Introduction.—The physics of quantum thermalization, which describes how a system reaches an equilibrium state, has garnered substantial attention in recent years [1, 2, 3, 4, 5, 6, 7, 8]. A key concept – the eigenstate thermalization hypothesis [9, 1, 10, 7] – indicates that closed systems with unbroken ergodicity invariably reach an equilibrium state at late times. However, not all systems behave this way, a notable exception being the observation of persistent oscillations in a 51 Rydberg atom quantum simulator [11]. This form of ergodicity breaking has been attributed to non-thermalizing eigenstates known as quantum many-body scars (QMBSs) [12, 13] (see also the reviews [14, 15, 16]). In contrast to integrable models [17] or strongly disordered systems displaying many-body localization [18, 19], typical QMBS systems weakly break ergodicity due to the non-thermal eigenstates forming a small-dimensional representation of a “restricted spectrum generating algebra” [20, 21, 22, 23, 24], although more general mechanisms continue to be explored [25, 26, 27, 28].
The experimental observations of QMBSs in versatile synthetic platforms ranging from ultracold atoms to superconducting circuits [11, 29, 30, 31, 32, 33, 34] have revealed the intrinsic fragility of QMBS dynamical revivals. In some cases, this has been mitigated by periodic driving, which can extend the lifetime of QMBS oscillations [29, 32, 35, 36]. Nevertheless, thermalization or decay to a steady state, due to inherent decoherence from atom loss or thermal fluctuations, ultimately sets in. Thus, achieving stable QMBSs in real-world settings, with unavoidable non-unitarity, remains a challenge for both theory and experiment [37, 38, 39, 40, 41].
In this paper, we show that non-Hermiticity presents a new route for enhancing ergodicity breaking and many-body scarring. Borrowing inspiration from well-known mechanisms such as non-Hermitian pumping [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55], we demonstrate that a class of many-body systems with kinetic constraints can admit a form of “Fock skin effect” that amplifies the revivals of special initial states and makes them resilient against spatial disorder. Unlike the conventional non-Hermitian skin effect (NHSE), the Fock skin effect is characterized not by the uni-directional spatial evolution of wavefunctions, but by persistent dynamical recurrences in the Fock space of a many-body system. We demonstrate this mechanism of QMBS enhancement in a non-Hermitian variant of the PXP model, which can be realized in a Bose–Hubbard optical lattice with laser-induced loss [56, 57, 58, 59, 60]. Connection with conventional non-Hermitian pumping is made through the “forward scattering approximation” [12, 61, 22], which effectively maps the many-body system to a spatially-inhomogeneous tight-binding chain with non-reciprocal hopping amplitudes. Our results offer a new perspective on obtaining stable many-body revivals through dissipative perturbations, extending beyond recent approaches based on decoherence-free subspaces [62, 63].
Figure 1: (a) The action of operators in Eq. (1). Blue directed arrows label the asymmetric couplings $\hat{P}_{j-1}\hat{X}^{\prime}_{j}\hat{P}_{j+1}$ (arrow for enhanced spin flipping $(1{+}u)\hat{X}^{+}_{j}$), while red links represent the symmetric coupling $\hat{P}_{j-1}\hat{X}_{j}\hat{P}_{j+1}$. (b) The directed Fock-space graph for the non-Hermitian PXP model in Eq. (1) with $L=6$ sites under PBCs. Each vertex represents a Fock state. The vertices are organized into vertical layers according to their Hamming distance $x$ from the state $|\ket{\mathbb{Z}_{2}}\equiv\ket{\uparrow\downarrow\uparrow\downarrow\ldots}$, located at the left edge of the graph. The translated Néel state, |¯$|\ket{\bar{\mathbb{Z}}_{2}}\equiv\ket{\downarrow\uparrow\downarrow\uparrow\ldots}$, is at the opposite end of the graph. The edges between Fock states represent the transitions allowed by the Hamiltonian in Eq. (1). Notably, all directed edges point towards $x{=}0$, denoting asymmetric pumping from the Fock skin effect.
Fock skin effect.—We consider the following non-Hermitian generalization of the one-dimensional (1D) PXP model [64, 65], a paradigmatic model of QMBSs [12, 13]:
$\hat{H}=\sum^{L/2-1}_{n=0}\hat{P}_{2n-1}\hat{X}^{\prime}_{2n}\hat{P}_{2n+1}+\hat{P}_{2n}\hat{X}_{2n+1}\hat{P}_{2n+2}.$ (1)
Here, $\hat{X}$, $\hat{Y}$, $\hat{Z}$ denote the standard Pauli matrices, such that $\hat{P}{=}(\hat{I}{-}\hat{Z})/2$ projects onto the spin-$\downarrow$ state. Non-Hermiticity enters through the asymmetric spin flip operator, $\hat{X}^{\prime}{=}\hat{X}{+}iu\hat{Y}$, acting on the even sites, FIG. 1(a), where $u$ controls the strength of non-Hermiticity. Each symmetric $\hat{X}$ or antisymmetric $\hat{X}^{\prime}$ spin flip acts only when its neighboring spins are both spin-$\downarrow$. To distinguish our Fock skin effect from the conventional skin effect marked by spatial boundary accumulation, we remove the physical boundaries by imposing periodic boundary conditions (PBCs). Moreover, we restrict to the $10{<}u{<}$ 1 regime, where the energy spectrum is real, circumventing wavefunction decay induced by complex spectra, see Supplementary Material (SM) [66].
To provide insight behind the non-Hermitian scar enhancement, we analyze the wavefunction dynamics within the Fock space. The latter can be visualized as a quasi-1D network structure by grouping Fock states according to their Hamming distance from the Néel state, $|\ket{\mathbb{Z}_{2}}=\ket{\uparrow\downarrow\uparrow\downarrow\cdots}$, see an example in FIG. 1. Note that the |$\ket{\mathbb{Z}_{2}}$ state is the one that exhibits QMBS revivals in the Hermitian PXP model [12]. The Hamming distance indicates how many applications of the Hamiltonian are needed to reach the states in a given layer starting from the |$\ket{\mathbb{Z}_{2}}$ state. This is because the Hamiltonian only connects layer $x$ with $x\pm 1$, as shown by edges in FIG. 1(b), making apparent the connection with a tight-binding chain. Crucially, this picture reveals a consistent direction in the asymmetry of the transitions between Fock states. As illustrated, the blue directional transitions are all directed towards the |$\ket{\mathbb{Z}_{2}}$ state at $x{=}0$, reminiscent of a conventional NHSE chain but with Fock states instead of physical sites. However, with multiple Fock states in each layer instead of a single site, the “boundary” skin localization would be weak [66], and we instead expect more pronounced QMBS revivals, rather than the directional bulk current in conventional NHSE systems [43, 46, 67, 68].
The stratification of the Fock space into Hamming layers can be formally expressed by generalizing the forward scattering approximation (FSA) [12, 61], which allowed to understand the origin of QMBSs in the Hermitian case [22]. We first decompose our non-Hermitian PXP Hamiltonian in Eq.(1) as $\hat{H}{=}\hat{H}^{+}{+}\hat{H}^{-}$, where
$\hat{H}^{+}=$ $\sum_{j=2n}\hat{P}_{j-1}\hat{X}^{\prime-}_{j}\hat{P}_{j+1}\!+\!\sum_{j=2n+1}\hat{P}_{j-1}\hat{X}^{+}_{j}\hat{P}_{j+1},$ (2)
$\hat{H}^{-}=$ $\sum_{j=2n}\hat{P}_{j-1}\hat{X}^{\prime+}_{j}\hat{P}_{j+1}\!+\!\sum_{j=2n+1}\hat{P}_{j-1}\hat{X}^{-}_{j}\hat{P}_{j+1},$
with $\hat{X}^{\pm}{\equiv}(\hat{X}\pm i\hat{Y})/2$ and $\hat{X}^{\prime\pm}{\equiv}(1\pm u)(\hat{X}\pm i\hat{Y})/2$. By successively acting with $\hat{H}^{+}$ on the |$\ket{\mathbb{Z}_{2}}$ state, which we label as |$\ket{0}_{\rm FSA}$, we can generate a tower of $L+1$ FSA basis states |$\ket{n}_{\mathrm{FSA}}{\propto}(H^{+})^{n}\ket{0}_{\mathrm{FSA}}$, with $n=0,1,2,\ldots,L$ [61]. These states span the Fock-space layers in FIG. 1(b). In the FSA basis, the projected Hamiltonian (1) is given by
$\hat{H}_{\rm FSA}\!\!=\!\!\sum_{x=0}^{L}t^{-}_{x}|x\rangle_{\rm FSA}\langle x+1|_{\rm FSA}+t^{+}_{x}|x+1\rangle_{\rm FSA}\langle x|_{\rm FSA},$ (3)
with the non-uniform and asymmetric hopping amplitudes
$t^{+}_{x}\!=\!\bra{x\!+\!1}_{\rm FSA}\!\!\hat{H}^{+}\!\!\ket{x}_{\rm FSA},\quad t^{-}_{x}\!=\!\bra{x}_{\rm FSA}\!\!\hat{H}^{-}\!\!\ket{x\!+\!1}_{\rm FSA}\!.$ (4)
Despite the formal analogy between the model (3) and a tight-binding chain, the FSA basis |$\ket{x}_{\mathrm{FSA}}$ represents the superposition of all Fock states spanning Hamming layer $x$. Moreover, both the basis states and the effective hopping amplitudes $t^{\pm}$ intricately depend on the non-hermiticity $u$ of the Hamiltonian, as we show below.
Figure 2: (a) Normalized fidelity $F(t)$ for the initial state |$\ket{\mathbb{Z}_{2}}$ in the non-Hermitian PXP model (1). Turning on moderate non-Hermiticity $u\neq 0$ significantly enhances the return probability to the initial state. For large $u$ (dashed green curve), robust revivals are still present, but they become irregular due to multiple frequencies in the dynamics. (b) Amplitudes of effective hoppings in Eq.(4). The asymmetry between $t^{\pm}_{n}$ for $u{>}0$ (circles vs. triangles) leads to the non-Hermitian Fock skin effect in our model. (c) The concentration of eigenstates in the Hamming layer $x$, Eq. (5). A sharp Fock skin mode emerges under strong non-Hermiticity (green curve), corresponding to the persistent revivals in (a). (d) Comparison between the FSA left-boundary localization, $|\bra{\mathbb{Z}_{2}}\ket{\phi_{j}}_{\rm FSA}|$ |, and overlaps of |$\ket{\mathbb{Z}_{2}}$ with all eigenstates of the Hamiltonian in Eq.(1). From $u{=}0.0$ to $u{=}0.2$, the boundary localization from FSA vectors is in close agreement with the overlap of $\mathbb{Z}_{2}$ state with the equally-spaced QMBS states. All data is for a chain of size $L{=}16$ with PBCs.
QMBS enhancement due to Fock skin effect.—QMBSs manifest as pronounced revivals in the fidelity, $F(t)\equiv|{\bra{\psi(0)}\ket{\psi(t)}}|^{2}/{\bra{\psi(t)}\ket{\psi(t)}}$, following a global quench from a chosen initial state, typically the |$\ket{\psi(0)}=\ket{\mathbb{Z}_{2}}$ state for PXP-type models [11, 12]. The time evolved state is obtained by |$\ket{\psi(t)}=\exp(-i\hat{H}t)\ket{\psi(0)}$, and we explicitly normalize the fidelity at each time step. Strong revivals in $F(t)$ indicate broken ergodicity since the fidelity should decay rapidly under thermalizing dynamics [12, 13]. In FIG. 2(a), the fidelity $F(t)$ is plotted for the PXP model (1) with various degrees of non-Hermiticity $u$. The decay of peaks in $F(t)$ is noticeably slower for non-Hermitian $u{=}0.2$ (blue) as compared to the Hermitian $u{=}0$ case (red). With much larger non-Hermiticity $u{=}0.7$ (green dashed), both frequency and magnitude of the revivals are significantly enhanced, such that no significant decay is observed even after $t{=}10$.
To explicitly attribute this non-Hermitian scar enhancement to the Fock skin effect, we plot the hopping amplitudes (4) in FIG. 2(b). Saliently, the FSA hoppings are asymmetric with $t_{x}^{+}$ (round markers) and $t_{x}^{-}$ (triangle markers) of different magnitudes, with the asymmetry becoming more pronounced with greater non-Hermiticity $u$. We note that the hopping asymmetry is markedly different from that in the physical Hamiltonian (1), since $x$ is the Hamming distance from the |$\ket{\mathbb{Z}_{2}}$ state in Fock space and not a spatial coordinate. Importantly, the asymmetric hopping amplitudes are non-trivially dependent on $x$ but asymmetric in the same direction ($t_{x}^{+}\geq t_{x}^{-}$) within the entire FSA chain. We emphasize that this property is not generic, see SM [66] for the discussion of a related model where this is not the case, with crucial implications on the existence of Fock skin accumulation.
The FSA non-uniformity and the large dimensionality of each layer $x$ lead to a unique Fock skin effect, distinct from typical skin localization. This is elucidated through the Hamming distance-resolved Fock space density $\Pi_{\rm Fock}(x)$ which quantifies the total weight of all eigenstates at a Hamming distance $x$ from the |$\ket{\mathbb{Z}_{2}}$ state:
$\Pi_{\rm Fock}(x)=\frac{1}{\mathcal{D}}\sum_{\ket{\varphi}\in\mathcal{L}(x)}\sum_{i}|\bra{\varphi}\ket{\phi_{i}}|^{2},$ (5)
with $\mathcal{L}(x)$ denoting the set of all Fock states in the Hamming layer $x$ [69], |$\ket{\phi_{i}}$ is the $i$-th eigenstate, and $\mathcal{D}$ is the dimension of the Hilbert space.
As seen in FIG. 2(c), the Fock skin pumping of $u{=}0.2$ shifts the symmetric $\Pi_{\rm Fock}(x)$ profile of $u{=}0$ slightly towards left, yet without pronounced “boundary” localization. Nevertheless, this increased proximity to the |$\ket{\mathbb{Z}_{2}}$ state is sufficient to noticeably enhance the scarred revivals [FIG. 2(a), blue]. A much stronger non-Hermiticity of $u{=}0.7$ results in a much more skewed $\Pi_{\rm Fock}(x)$, but still much less localized than conventional skin states corresponding to $t_{x}^{+}/t_{x}^{-}{\approx}3$ [FIG. 2 (b), green], with localization lengths of about $[\log 3]^{-1}{\approx}1$ site. Crucially, this lack of conventional exponential boundary skin localization allows for significant state diffusion in the Hilbert space, but with enhanced return probability to the |$\ket{\mathbb{Z}_{2}}$ state.
We next explore how shifted Fock skin profiles amplify QMBS revivals. Within the FSA framework, FIG. 1, scarred revivals oscillate from the left FSA boundary |$\ket{0}_{\rm FSA}\equiv\ket{\mathbb{Z}_{2}}$ to the opposite end, |$\ket{L}_{\rm FSA}\equiv\ket{\bar{\mathbb{Z}}_{2}}$. The revivals are driven by a branch of eigenstates |$\ket{\phi}$ with approximately equally spaced eigenenergies $E$ and high overlaps $|\bra{\mathbb{Z}_{2}}\ket{\phi}_{\rm FSA}|$ |, as depicted in FIG. 2(d). These FSA overlaps (hollow markers) closely align with the overlaps of |$\ket{\mathbb{Z}_{2}}$ with the eigenstates of our original PXP Hamiltonian (1) (solid markers), demonstrating the accuracy of the FSA description. In particular, the Fock skin effect boosts the |$\ket{\mathbb{Z}_{2}}$ overlap of the FSA branch at $u{=}0.2$, consequently enhancing revivals over the Hermitian case at $u{=}0.0$. Greater non-Hermiticity, e.g., $u{=}0.7$, enhances the |$\ket{\mathbb{Z}_{2}}$ overlap even more but also introduces other non-thermal states with high overlaps, such that additional frequencies appear in the revival dynamics – see the green curve in FIG. 2(a).
Robustness to disorder.—We have established that the Fock skin effect can significantly enhance QMBSs. Given that the real-space skin effect can result in robust, directed dynamical currents, we posit that the non-Hermiticity in our system can not only amplify the quantum scar but also protect it against external (Hermitian) perturbations. To check this, we investigate the robustness of quantum scar under spatial disorder,
$\hat{H}_{\rm dis}=\hat{H}+\sum_{j}w_{j}\hat{Z}_{j},$ (6)
where $w_{j}$ denotes the strength of on-site random potential drawn from a uniform distribution within the range $[-W/2,W/2]$. It was previously shown that Hermitian QMBS revivals decay with increasing disorder strength [70]. For our model with the Fock skin effect at up to $u{=}0.35$, this resilience under moderate disorders is remarkably highlighted by the fidelity shown in FIG. 3(a), where robust scarring signatures are still present, in contrast to the rapid collapse at Hermitian $u{=}0$ without the Fock skin effect. The quantitative demonstration of enhanced stability is illustrated through the first fidelity peak, $F_{1}={\rm Max}[F(t)]$ with $t\in[2.5,7.5]$, as shown in FIG. 3 (b). The peaks are markedly enhanced with increasing $u$, and the subsequent enhanced revivals will be significantly more pronounced.
Figure 3: (a) The representative fidelity dynamics for the disordered Hamiltonian $\hat{H}_{\rm dis}$ in Eq. (6) at several values of $u$. At $u{=}0$, the scarring signatures swiftly collapse at moderate disorders. The Fock skin effect at up to $u{=}0.35$ leads to prolonged and robust scarring recurrences. (b) The first fidelity peak $F_{1}$, observed with $t\in[2.5,7.5]$, which is significantly shifted by non-Hermiticity $u$. These results in (b) are based on $200$ realizations, and the error bar denotes the standard deviation obtained throught such $200$ realizations. All the results in (a) and (b) are computed under conditions: system size $L{=}14$, |$\ket{\mathbb{Z}_{2}}$ initial state, and uniform disorder strength of $W=0.8$.
Optical lattice proposal.—Finally, we outline a proposal for observing non-Hermitian QMBSs in existing Bose-Hubbard (BH) quantum simulators with additional laser-induced loss [71, 72, 73, 32], see FIG. 4. Such a simulator is described by following BH-type Hamiltonian (see SM [66] for a derivation)
$\hat{H}_{\mathrm{BH}}\!\!$ $=\!\!-\!\!\!\!\!\sum_{j=0}^{L/2-1}\!\![\!\left.\left(J\!\!+\!\tilde{\gamma}\right)\hat{b}_{2j}^{\dagger}\hat{b}_{2j\!+\!1}\!+\!\!\right.\left(J\!\!-\!\tilde{\gamma}\right)\hat{b}_{2j\!+\!1}^{\dagger}\hat{b}_{2j}\!\!+\!\!J\hat{b}_{2j\!+\!2}^{\dagger}\hat{b}_{2j\!+\!1}\!\!$ (7)
$+J\hat{b}_{2j+1}^{\dagger}\hat{b}_{2j+2}]+\!\!\!\sum_{j=0}^{L-1}[\Delta j\hat{n}_{j}\!\!+\!\!\frac{U}{2}\hat{n}_{j}\left(\hat{n}_{j}\!-\!1\right)],$
where $\hat{b}_{j}^{\dagger}$ ($\hat{b}_{j}$) denotes the boson creation (annihilation) operator on site $j$, and $\hat{n}_{j}{=}\hat{b}^{\dagger}_{j}\hat{b}_{j}$. The nonreciprocity $J{\pm}\tilde{\gamma}$ results from the adiabatic elimination of a dissipative excited state (wavy arrow in FIG. 4 for laser-induced atom loss) linked to the nearest ground states [59]. The potential tilt $\Delta$ and on-site repulsion $U$ induce kinetic constraints. The effective non-Hermitian PXP model (1) arises when we set the boson filling factor to $\nu{=}N/L{=}1$ and $U{=}\Delta\gg J,\gamma$ [66]. Under these conditions, the local transition between the “20” doublon and the “11” neighboring single-boson configurations can be exactly mapped to the Pauli spin flip operator $\hat{X}$ [74, 75, 32, 66], as shown by the red box in Fig. 4 (a), such that the initial |$\ket{\mathbb{Z}_{2}}$ state corresponds to a product state of doublons separated by empty sites, i.e. “$2020$”.
Enhanced scarring can be experimentally observed by measuring the total imbalance of the boson density across odd and even sites, $I(t){=}(1/N){\sum_{i}\left(n_{2i}(t){-}n_{2i+1}(t)\right)}$, where $n_{i}(t){=}\bra{\psi(t)}\hat{n}_{i}\ket{\psi(t)}/\bra{\psi(t)}\ket{\psi(t)}$ is the normalized occupation of site $i$ and $N$ is the total number of bosons. $I(t)$ is proportional to the Néel order parameter or |$\ket{\mathbb{Z}_{2}}$ echo corresponding to the effective spins. Simulation results are shown in FIG. 4(b) with $V{=}\Delta{=}10J$. Saliently, the non-Hermitian case with $\tilde{\gamma}{=}0.3J$ (red) exhibits stronger revivals of $I(t)$ than the Hermitian case with vanishing $\tilde{\gamma}$ (blue).
Figure 4: (a) Schematics of the non-Hermitian Bose-Hubbard model (7) in a tilted optical lattice. The non-reciprocal hopping $J\pm\tilde{\gamma}$ (black arrow) is realized by coupling (blue and violet double arrows) two nearest ground states with an excited state with the laser-induced loss (wavy arrow) on the excited state. $U$ represents the on-site interaction strength, while $\Delta$ refers to the linear potential induced by tilting the optical lattice. Kinetic constraints emerge under $U{=}\Delta\gg J,\tilde{\gamma}$ [32], resulting in the effective non-Hermitian PXP model in Eq.(1). This correspondence is illustrated in the red block that the boson hopping is mapped to spin flipping [66]. (b) Dynamics of population imbalance $I(t)$ for the non-Hermitian Bose-Hubbard model (7), as an experimental indicator for enhanced quantum scars. The initial state is ``$20202020201"$ with $L{=}11$. Other parameters are $V{=}\Delta{=}10J$, and $\tilde{\gamma}/J{=}0.0$ (blue) and $0.3$ (red). The red curve exhibits pronounced and robust revivals, showing the enhancement of quantum scars.
Discussion.—We have shown that asymmetric Fock-state transitions can significantly enhance quantum many-body scarring. Unlike the conventional NHSE, which appears at the physical boundaries of a system, our Fock skin effect leads to many-body state revivals through a delocalized Fock skin accumulation. The latter are enabled by consistent asymmetric transitions across the Hamming layers associated with a particular initial state. We have demonstrated that the dynamics of non-Hermitian scar enhancement can be realized in existing Bose-Hubbard quantum simulators [73, 32]. Furthermore, the rapidly improving quantum computing hardware [76, 77, 78, 79, 80], combined with advancements in quantum algorithms, such as variational circuit optimization [81, 82, 83, 84], non-unitary postselection [85, 81, 83] and advanced error mitigation methods [86, 87], provides another platform for simulating our proposed non-Hermitian PXP and related many-body models  [88, 82, 89, 90].
The mechanisms for stabilizing scar states in disordered systems have recently attracted much attention [91, 92, 93, 94, 95, 96, 97, 98]. The robustness of QMBSs in our model is protected by the Fock skin effect, which is absent in other types of non-Hermitian deformations of the PXP model [99], see SM [66]. Moreover, compared to previous works on disordered scared systems [92], the non-Hermiticity in our case can “shield” all characteristic QMBS signatures, including wavefunction revivals, from disorder. Such robust revivals could also be useful for realizing stable dissipative time crystals under various perturbations [100, 35]. Finally, it would be interesting to explore the consequences of the Fock skin effect introduced here for entanglement phase transitions [101, 102] in various non-Hermitian many-body systems [103, 104, 105, 106, 107, 83, 108].
Acknowledgements.
Acknowledgements.—Exact diagonalization simulations have been performed using QuSpin [109, 110]. F.Q. and C.H.L acknowledges support from the QEP2.0 Grant from the Singapore National Research Foundation (Grant No. NRF2021-QEP2-02-P09) and the Singapore MOE Tier-II Grant (Proposal ID: T2EP50222-0008). J.-Y.D. and Z.P. acknowledge support by the Leverhulme Trust Research Leadership Award RL-2019-015. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No.101034413. This research was supported in part by grants NSF PHY-1748958 and PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
Supplemental Online Material for “Enhanced many-body scars from the non-Hermitian Fock space skin effect”
[…]
${}^{1}$epartment of Physics, National University of Singapore, Singapore 117542
${}^{2}$chool of Physics and Astronomy, University of Leeds, Leeds LS2 9JT, United Kingdom
${}^{3}$nstitute of Science and Technology Austria (ISTA), Am Campus 1, 3400 Klosterneuburg, Austria
This supplementary material contains: (SI) a detailed description of the experimental setup for realizing an effective non-Hermitian PXP model using ultracold atoms in an optical lattice; (SII) further characterization of the Fock skin accumulation in three variations of the model with different types of non-Hermiticity; (SIII) analysis of the energy spectrum as a function of non-Hermiticity; (SIV) additional evidence for the enhancement of scarring by the non-Hermiticity in the fidelity dynamics and entanglement entropy of eigenstates.
Appendix A SI. Effective non-Hermitian PXP model from a non-Hermitian Bose-Hubbard model
In the main text, we introduced a non-Hermitian generalization of the PXP model and showed that its non-Hermiticity enhances the quantum-many body scars (QMBSs). Here we show how that model can effectively emerge in a Bose-Hubbard quantum simulator in the presence of laser-induced loss. Below we provide details of our proposed experimental setup and measurement, building upon previous work in the Hermitian setting [32].
We consider ${}^{87}$b atoms in a one-dimensional (1D) optical lattice, FIG. S1(a), described by a Bose-Hubbard type Hamiltonian defined in terms of standard boson annihilation and creation operators, $\hat{b}_{j}$ and $\hat{b}_{j}^{\dagger}$:
$\hat{H}_{\rm BH}=\hat{H}_{\rm hopping}+\sum_{j=0}^{L-1}\Delta j\hat{n}_{j}+\frac{U}{2}\sum_{j=0}^{L-1}\hat{n_{j}}\left(\hat{n}_{j}-1\right),$ (S1)
where $\hat{n}_{j}=\hat{b}^{\dagger}_{j}\hat{b}_{j}$ is the local boson number on site $j$. The three terms in the Hamiltonian stand for, respectively, the kinetic energy due to bosons hopping between lattice sites, the linear tilt potential of the optical lattice parametrized by $\Delta$, and the on-site interaction strength $U$ that penalizes the multiple occupancy of any site.
The hopping term in the BH model in Eq. (S1) is of non-Hermitian form, reminsiscent of the Su-Schrieffer-Heeger model:
$\hat{H}_{\mathrm{hopping}}$ $=-\sum_{j}\left\{\left(J+\tilde{\gamma}\right)\hat{b}_{2j}^{\dagger}\hat{b}_{2j+1}+\left(J-\tilde{\gamma}\right)\hat{b}_{2j+1}^{\dagger}\hat{b}_{2j}+J\hat{b}_{2j+2}^{\dagger}\hat{b}_{2j+1}+J\hat{b}_{2j+1}^{\dagger}\hat{b}_{2j+2}\right\}+\sum_{j}\tilde{\gamma}\hat{b}_{j}^{\dagger}\hat{b}_{j},$ (S2)
where $J$ is the hopping amplitude and the asymmetry characterized by $\bar{\gamma}$ is induced by atom loss, see FIG S1. The non-Hermitian asymmetric hoppings of strength $J\pm\tilde{\gamma}$ can be engineered as follows [58, 59]. For each unit cell shown in FIG. S1(a), we couple two nearest ground-state levels (denoted by $A$ and $B$) via an excited state ($e$ state) with Rabi frequency $\Omega_{0}$ (dashed arrows). The excitation towards state $e$ serves as atomic loss in the ground state manifold, which we quantify phenomenologically as $\gamma$. As sketched in FIG. S1(b), this leads to effective non-reciprocal hoppings in Eq.(S2) between atoms on the nearest ground states, which can be rigorously shown by adiabatically eliminating the excited state (see details in Ref. [59]). The strength of $\tilde{\gamma}$ in Eq. (S3) is given in terms of the Rabi frequency $\Omega_{0}$ by
$\tilde{\gamma}=\frac{\Omega_{0}^{2}}{\gamma+i\delta}\approx\frac{\Omega_{0}^{2}}{\gamma},$ (S3)
where $\delta$ is the detuning on the excited state and we assumed $\gamma\gg\delta$ [59].
In experiments, dynamics is conveniently described by measuring the total occupation imbalance between odd and even sites:
$I(t)=\frac{\sum_{i}\left(n_{2i}(t)-n_{2i+1}(t)\right)}{\sum_{i}\left(n_{2i}(t)+n_{2i+1}(t)\right)},$ (S4)
where $n_{j}=\bra{\psi(t)}\hat{b}_{j}^{\dagger}\hat{b}_{j}\ket{\psi(t)}$ is the occupation density on site $j$. For this quantity, we can drop the global loss $\sum_{j}{\gamma}\hat{b}_{j}^{\dagger}\hat{b}_{j}$ in Eq.(S2[111]. This is justified as any loss that affects all sites similarly does not impact the imbalance.
Figure S1: (a) Schematic of the non-Hermitian Bose-Hubbard model in Eq. (S3). The atoms are trapped in a tilted optical lattice with an alternating chain of local ground states “A” and “B” (yellow and blue dots). $U$ is the strength of the on-site interaction, $\gamma$ is the laser-induced loss rate on the excited state “e”, and $\delta$ denotes the strength of detuning on the excited state [58, 59]. (b) Adiabatic elimination of the excited state “e” (red box in (a)) results in the model in Eq. (S2[59]. Effective non-reciprocal hoppings $J\pm\tilde{\gamma}$ are realized by coupling the two nearest ground states “A” and “B” via an excited state “e” with Rabi frequency $\Omega_{0}$ (black dashed arrows), with $\tilde{\gamma}=\Omega_{0}^{2}/(\gamma+i\delta)$.
A.1 Effective non-Hermitian PXP model
To induce kinetic constraints in the Bose-Hubbard model in Eq. (S1), we tune the system to a resonance $U=\Delta\gg J,\tilde{\gamma}$ and assume the total number of bosons is equal to the number of lattice sites, corresponding to filling factor $\nu=N/L=1$. In the absence of atomic loss ($\tilde{\gamma}=0$), Ref. [32] observed that under this resonance condition, the two-body repulsion $U$ and linear potential $\Delta$ interplay in such a way that any triple occupancy on a single site is heavily suppressed, leaving only the transitions of the type“$11$$\leftrightarrow$20202020”, i.e., between two singly occupied adjacent sites and a doubly occupied site. This allows to map the dynamics onto a simpler PXP spin model. Below we will generalize this to the case where atomic loss is present. In terms of experimental implementation, the effective model derived below is expected to hold in the regime $U=\Delta=10J\approx 2.0$ kHz. The tilt potential $\Delta$ can be realized either with gravity or external magnetic fields; for instance, in the setup of Ref. [32] $\Delta_{g}=1632$ Hz from gravity (tilting degree $4^{\circ}$ and $767$ nm spacing for ${}^{87}$b atoms), and $\Delta_{B}=368$ Hz from external magnetic fields, resulting in the quoted value $\Delta=\Delta_{g}+\Delta_{B}\approx 2$ kHz.
Working at the resonance $U=\Delta\gg J,\tilde{\gamma}$, the dynamics governed by the model in Eq. S1 can be describde by the following effective Hamiltonian [32]
$\hat{H}_{\mathrm{eff}}=$ $-\sum_{j=2n}(\left(J+\tilde{\gamma}\right)\underbrace{\hat{b}_{j}^{\dagger}\hat{b}_{j+1}\delta_{\hat{n}_{j},1}\delta_{\hat{n}_{j+1},1}}_{\sqrt{2}\hat{P}_{j-1}\hat{X}_{j}^{+}\hat{P}_{j+1}}+\left(J-\tilde{\gamma}\right)\underbrace{\hat{b}_{j+1}^{\dagger}\hat{b}_{j}\delta_{\hat{n}_{j},2}\delta_{\hat{n}_{j+1},0}}_{\sqrt{2}\hat{P}_{j-1}\hat{X}_{j}^{-}\hat{P}_{j+1}})$ (S5)
$-\sum_{j=2n+1}(J\underbrace{\hat{b}_{j}^{\dagger}\hat{b}_{j+1}\delta_{\hat{n}_{j},1}\delta_{\hat{n}_{j+1},1}}_{\sqrt{2}\hat{P}_{j-1}\hat{X}_{j}^{+}\hat{P}_{j+1}}+J\underbrace{\hat{b}_{j+1}^{\dagger}\hat{b}_{j}\delta_{\hat{n}_{j},2}\delta_{\hat{n}_{j+1},0}}_{\sqrt{2}\hat{P}_{j-1}\hat{X}_{j}^{-}\hat{P}_{j+1}}).$
As indicated below each term in the effective Hamiltonian, different processes have a simple spin-1/2 representation with the Fock states mapped to spin product states as follows:
[…]
$\hat{X}^{+}=\left[\begin{array}[]{ll}0&1\\ 0&0\end{array}\right],\;\;\;\hat{X}^{-}=\left[\begin{array}[]{ll}0&0\\ 1&0\end{array}\right].$ (S7)
Finally, the effective Hamiltonian in Eq. (S5) can be written as the following non-Hermitian PXP Hamiltonian
$\hat{H}_{\mathrm{NHPXP}}^{\prime}=\Omega\sum_{j=2n}\hat{P}_{j-1}\hat{X}^{\prime}_{j}\hat{P}_{j+1}+\Omega\sum_{j=2n+1}\hat{P}_{j-1}\hat{X}_{j}\hat{P}_{j+1}$ (S8)
where $\hat{P}=(1-\hat{Z})/2$, and we have introduced
$\hat{X}=\left[\begin{array}[]{ll}0&1\\ 1&0\end{array}\right],$ (S9)
$\hat{X}^{\prime}=\left[\begin{array}[]{cc}0&1+u\\ 1-u&0\end{array}\right]=\hat{X}+iu\hat{Y},$ (S10)
$\Omega=\sqrt{2}J,\;\;\;u=\tilde{\gamma}/J.$ (S11)
For the quench dynamics, we are primarily interested in the $\left|\mathbb{Z}_{2}\right\rangle=\left|\uparrow\downarrow\uparrow\downarrow\uparrow\downarrow\uparrow\downarrow\uparrow\downarrow......\right\rangle$ product state of spins, which corresponds to the product state of doublons, $\left|202020......\right\rangle$, in the Bose-Hubbard model. The latter can be experimentally prepared using an optical superlattice [32].
Appendix B SII. Fock skin accumulation and hopping inhomogeneity in the FSA chain
As discussed in the main text, the origin of the enhanced scar in our model is the non-Hermitian skin accumulation occurring within the Fock space, which can be understood as the accumulation in the FSA chain. Contrary to the uniform non-reciprocity of the physical hoppings, the hopping asymmetry in the effective FSA chain is highly non-uniform, potentially even flipping at different parts of the chain. This is because the chain is along the Hamming distance $x$ from the |$\ket{\mathbb{Z}_{2}}$ state, and each “site” actually represents an extensively large number of Fock basis states.
To better understand the nature of the Fock skin effect in our model, Eq. (S8), we compare three different variations of it, labeled by $a$, $b$, and $c$:
$\hat{H}_{\mathrm{a}}=\sum_{j}\hat{P}_{j-1}\hat{X}^{\prime}_{j}\hat{P}_{j+1},$ (S12)
${\bf Primary~{}model:}\hat{H}_{b}=\sum_{j=2n}\hat{P}_{j-1}\hat{X}^{\prime}_{j}\hat{P}_{j+1}+\sum_{j=2n+1}\hat{P}_{j-1}\hat{X}_{j}\hat{P}_{j+1},$ (S13)
$\hat{H}_{c}=\sum_{j=2n}\hat{P}_{j-1}\hat{X}^{\prime}_{j}\hat{P}_{j+1}+\sum_{j=2n+1}\hat{P}_{j-1}(\hat{X}^{\prime}_{j})^{\dagger}\hat{P}_{j+1}.$ (S14)
Note that this model $b$ describes our experimental setup in Eq. S8 with unit $\Omega=1$.
The Hamiltonian $\hat{H}_{a}$ describes the model considered in Ref. [99], which is invariant under translation. Model $b$ is the one we investigate in the main text, as $\hat{H}_{b}$ is equivalent to Eq. (S8) with $\Omega=1$. Finally, we introduce the model $c$ as a more extreme version of the model $b$. Indeed, in the model $c$ the non-Hermitian elements are staggered to mimic the structure of the Néel state. This means that, in the graph representation of FIG. 1 in the main text, not only are all couplings asymmetric (unlike in our model $b$) but they all point towards the Néel state (unlike in model $a$). This is expected to make the scarring more prominently linked to the Fock skin effect. However, the presence of $(\hat{X}^{\prime}_{j})^{\dagger}$ in $\hat{H}_{c}$ makes it much more complicated to implement in a Bose-Hubbard quantum simulator. For that reason, in the main text, we only consider model $b$.
Figure S2: Fock skin accumulation in various models, Eqs. (S12)-(S14). All the panels show the Fock-space density, defined by Eq. (S15), for the system size $L=12$. In the Hermitian case $u=0$, $\Pi_{\mathrm{Fock}}(x)$ is perfectly symmetric and there is no Fock skin accumulation. For the model $a$, there is still no accumulation even with non-zero $u$. On the other hand, as expected, models $b$ and $c$ show a strong bias towards $x=0$ when $u>0$.
In the main text, we have introduced the density in Fock space
$\Pi_{\rm Fock}(x)=\frac{1}{\mathcal{D}}\sum_{i}\sum_{\ket{\varphi}\in\mathcal{L}(x)}|\bra{\varphi}\ket{\phi_{i}}|^{2},$ (S15)
where $\mathcal{L}(x)$ denotes the set of states in the Hamming layer $x$ [69], |$\ket{\phi_{i}}$ is the $i$th eigenstate, and $\mathcal{D}$ is the total number of Fock states. The Fock-space density for the above three models is shown in FIG. S2. For our proposed improvement of QMBS through the Fock skin effect, a key criterion is that the non-Hermiticity must generate an overall asymmetric flow in the entire Fock space with respect to the scarred trajectory. However, here we find that the Fock skin modes only appear in models $b$ and $c$, as revealed by the asymmetric profiles in FIGs. S2(b)-(c).
To understand why no Fock skin accumulation appears in the model $a$, we examine the symmetrized transition amplitudes between Hamming layers, as given by
$T^{+}_{x}=\bra{\tilde{x}}H\ket{\tilde{x}+1}\quad\text{and}\quad T^{-}_{x}=\bra{\tilde{x}+1}H\ket{\tilde{x}},$ (S16)
where |$\ket{\tilde{x}}=\sum_{\ket{\varphi}\in\mathcal{L}(x)}\ket{\varphi}$ is a symmetric superposition of all the states in the Hamming layer $x$. It is important to note that $T^{\pm}_{x}$ differ from the $t^{\pm}_{x}$ FSA hoppings used in the main text. While $t^{\pm}_{x}$ are meant to capture the scarred dynamics, $T^{\pm}_{x}$ give a more general idea of a state flow between the Hamming layers. Note that $T^{\pm}_{x}$ only meaningfully represent the individual physical Fock state transitions when $-1\leq u\leq 1$, where all non-zero matrix elements of the physical Hamiltonian are positive in the Fock basis, and hence there can be no cancellation between different terms.
Figure S3: Symmetrized transition amplitudes $T^{\pm}_{x}$, Eq. (S16), between the Hamming layers for the models in Eqs. (S12)-(S14). The system size is $L=12$ in all the plots. Note that for the model $a$, $T^{+}_{x}$ is larger than $T^{-}_{x}$ on the right, and vice versa on the left, hence leading to no consistent Fock skin pumping.
To have the Fock skin effect for scar enhancement, instead of merely strong boundary localization, two conditions must be met: (i) consistent asymmetry of the transitions $T^{-}_{x}>T^{+}_{x}$, and (ii) $T^{\pm}_{0}\rightarrow 0$ which prevents strong skin accumulation at the $x=0$ boundary. The latter condition is necessary to prevent outright exponential skin accumulation at $x=0$, which will largely prevent the dynamical state from exploring the other Hamming layers. In light of these conditions, we first examine the hopping amplitudes of Fock-space Hamiltonians in FIG. S3. The model $a$ in Eq. (S12) fails to meet the condition $T^{-}_{x}>T^{+}_{x}$, as it has opposite asymmetry in the left/right parts of the chain. This accounts for the previous lack of any Fock-space NHSE in FIG. S2(a). Conversely, for models $b$ and $c$, the non-reciprocity, as demonstrated by blue and green curves in FIGs. S3(b)-(c), generates an overall non-Hermitian pumping in the Fock space. Indeed, the model $c$, with greater $T^{\pm}_{x}$ hopping asymmetry at some $x$, gives rise to a very slightly stronger Fock skin accumulation than model $b$. Importantly, the condition $T^{\pm}_{0}\rightarrow 0$ ensures that there is no robust boundary localization.
As discussed in the main text, the resilience to perturbations stands out as a hallmark of the Fock skin effect. Thus, we examine the behavior of our models $a$, $b$ and $c$, subject to spatial disorder:
$\hat{H}_{\rm def}=\hat{H}+\sum_{j}w_{j}\hat{Z}_{j},$ (S17)
where $w_{j}\in[-W/2,W/2]$ is the strength of an on-site random potential [70]. As shown in FIG. S4 (a), models $b$ and $c$ exhibit relatively robust QMBS revivals even at strong disorder. This is attributed to the Fock skin effect. In contrast, model $a$ shows a rapid decay, akin to the behavior observed in the unitary scenario shown in the main text. Moreover, the first fidelity peak $F_{1}$ across system sizes $L$, as depicted in FIG. S4 (b), reveals that models b and c exhibit slower decay rates, compared to model a. Importantly, this shows the robustness of our enhanced scars, which are expected to be pronounced in physical experiments involving larger system sizes.
Figure S4: (a) Fidelity dynamics for the models in Eqs. (S12)-(S14), under the disorder of $W=1.8$ in Eq. (S17). We set $L=14$, $u=0.35$ and the |$\ket{\mathbb{Z}_{2}}$ initial state. Due to the absence of the Fock skin effect, the QMBS revivals for the model $a$ (blue curve) rapidly collapse. (b) The scaling behavior of the first fidelity peak $F_{1}$ within $t\in[2.5,7.5]$, obtained by $200$ realizations. We set $W=0.8$ and $u=0.4$ for (b).
To quantify the extent of Fock skin accumulation, and also to clarify that there is no boundary localization compared to the conventional NHSE, we define the mean position of the Fock-space density as follows
$\Pi^{\prime}_{\rm Fock}=\frac{1}{L}\sum_{x}x\Pi_{\mathrm{Fock}}(x).$ (S18)
This quantity reveals which Hamming layers the state mostly resides in. As illustrated in FIG. S2, a key factor linked to enhanced scar dynamics is the leftward shift of this density, due to the Fock skin effect. This leads to a shift of $\Pi^{\prime}$ towards $x=0$ for the models in Eq. (S13) and Eq. (S14), as shown in FIG. S5. However, importantly $\Pi^{\prime}_{\rm Fock}$ does not tend towards zero as $u$ increases to $0.6$. This is further supported by the effective hopping depicted in FIG. S3.
Figure S5: The center of Fock-space density, as defined in Eq. (S18) for $L=12$, compared across three models in Eqs. (S12)-(S14). Due to the absence of the Fock skin effect, the model $a$ does not exhibit a leftward shift. For models $b$ and $c$, the mean distance does not reach zero, indicating a lack of strong localization at the boundary of the Fock space.
Appendix C SIII. Real-complex transition in the spectrum of the non-Hermitian PXP model
Our PXP Hamiltonian can be expressed as follows
$\hat{H}=\sum_{i=1}^{N}\mathcal{P}\hat{X}_{i}\mathcal{P},$ (S19)
where the global projector is $\mathcal{P}=\prod_{i}\left[1-\left(1+\hat{Z}_{i}\right)\left(1+\hat{Z}_{i+1}\right)/4\right]$. Applying this global projector on the full Hilbert space, we can extract effective vectors, which lead to the following form
$H^{P}=\sum_{i=1}^{N}\hat{X}_{i},$ (S20)
Similarly, we can also express our non-Hermitian PXP in this sector
$\hat{H}_{\mathrm{NHPXP}}^{P}=\sum_{j=2n}\hat{X}^{\prime}_{j}+\sum_{j=2n+1}\hat{X}_{j}.$ (S21)
As such, the $\mathcal{P}\mathcal{T}$ symmetry breaking in our non-Hermitian setup appears at $u=1$ with the emergence of complex spectra, see FIG. S6 [112, 113, 114, 115, 116, 117, 118, 119, 120, 121]. Given the following form of $\hat{X}^{\prime}$
$\hat{X}^{\prime}=\left[\begin{array}[]{cc}0&1+u\\ 1-u&0\end{array}\right],$ (S22)
its singularity appears at $u=1$, and complex eigenvalues emerge under $u>1$. In the main text, we thus only consider the case $u<1$, where the spectrum of our non-Hermitian PXP model is purely real.
Figure S6: The real-complex transition at $u=1$ in our non-Hermitian PXP model in Eq.(S13) under PBCs.
Appendix D SIV. Entanglement entropy
In this section, we study the entanglement entropy of eigenstates of our non-Hermitian PXP model. As the focus of our work is state dynamics [101, 122] (of the $\mathbb{Z}_{2}$ Néel state), we will consider the entanglement of the right eigenvectors, rather than the biorthogonal entanglement entropy [123, 118, 124, 125, 126, 127].
The evolution of a wavefunction can be expressed using only the right eigenstates $\left|R_{i}\right\rangle$ with eigenvalues $E_{i}$ as
|$\ket{\psi(t)}=\sum_{i}C_{i}e^{-itE_{i}}\ket{R_{i}}.$ (S23)
Moreover, there is no significant skin effect along physical spaces. Thus, to characterize such dynamics through entanglement entropy, here we express the density matrix of our non-Hermitan model in terms of right eigenstates. We consider the half-chain entanglement entropy given by
$S\left(\rho^{RR}\right)=-\operatorname{Tr}[\rho^{RR}\log\rho^{RR}],$ (S24)
where $\rho^{RR}$ is the reduced density matrix of the left-half chain. The half-chain entanglement entropy $S$ is presented in FIG. S7. This allows us to identify and further characterize the scar eigenstates. Due to their large projection on a single product state |$\ket{\mathbb{Z}_{2}}$, the entanglement entropy of scarred eigenstates is expected to be significantly lower compared to typical thermalizing eigenstates at the same energy density. This is confirmed in FIG. S7 which shows the entanglement entropy of scarred eigenstates to be smaller and more separated from that of the thermal states for $u=0.2$ (blue) compared to $u=0$ (red), as highlighted by the dashed boxes. As these marked scar eigenstates are the main states underpinning the dynamics studied in the main text, the entanglement suppression can be related to the enhancement in the scarring dynamics.
Figure S7: Half-chain entanglement entropy in Eq. (S13) . We set $u=0.0$ for red dots and $u=0.2$ for blue dots. Note the significantly separated branch of low $S$ for the non-Hermitian case $u=0.2$ (blue). The system size is $L=16$.
Appendix E Structure of scarred revivals
To further characterize the QMBS revivals discussed in the main text, we plot the following echoes in FIG. S8
$E(t)=|\bra{\phi}\ket{\psi(t)}|^{2}/{\bra{\psi(t)}\ket{\psi(t)}},$ (S25)
with $\bra{\psi(t)}=\exp(-i\hat{H}t)\ket{\psi(0)}$, states |$\ket{\phi}=\ket{\mathbb{Z}_{2}}$ and |$\ket{\phi}=\ket{\mathbb{Z}^{\prime}_{2}}=\prod_{i}\hat{X}_{i}\ket{\mathbb{Z}_{2}}$. For $u=0.0$, we find that many-body revivals act as an oscillation between $\ket{\mathbb{Z}}$ and |$\ket{\mathbb{Z}^{\prime}_{2}}$, with the fidelity decaying with each state in-step. Meanwhile, for the non-Hermitian case $u=0.2$, these oscillations become unbalanced, with echoes to |$\ket{\mathbb{Z}_{2}}$ being stronger than for the Hermitian $u=0.0$ case. This is further exemplified for $u=0.5$ and $u=0.8$, where the dynamics and revivals predominantly occur only for the |$\ket{\mathbb{Z}_{2}}$ state. This leads to the more complex oscillations observed in our model. Nonetheless, we emphasize that the lifetime of the echoes to the |$\ket{\mathbb{Z}_{2}}$ state are already greatly improved for $u=0.2$, where the dynamics still explore most of the Hilbert space. This means that the improvement is not merely due to the wave function localizing around the |$\ket{\mathbb{Z}_{2}}$ state.
Figure S8: Echoes in Eq. (S25), with |$\ket{\phi}=\ket{\mathbb{Z}_{2}}$ for red curves and |$\ket{\phi}=\ket{\mathbb{Z}^{\prime}_{2}}=\prod_{i}\hat{X}_{i}\ket{\mathbb{Z}_{2}}$ for blue curves. For the Hermitian case with $u=0.0$, the dynamics exhibits balanced oscillation between states |$\ket{\mathbb{Z}_{2}}$ and |$\ket{\mathbb{Z}^{\prime}_{2}}$. Under the weak non-Hermiticity with $u=0.2$, the echo from |$\ket{\mathbb{Z}_{2}}$ state exhibits higher peaks, showing enhanced quantum scarring in our model. For the cases under strong non-Hermiticity under $u=0.5$ and $u=0.8$, more complicated oscillations emerge, and the echo of |$\ket{\mathbb{Z}^{\prime}_{2}}$ vanishes. System size is $L=14$ for all panels.